# Market-based insurance ratemaking: application to pet insurance

### v1

@@ 3.3 Population Monte Carlo Approximate Bayesian Computation algorithm
We start by setting a prior distribution $\pi(\theta)$ over the parameter space that we sequentially improve through intermediate distributions characterized by a sequence of tolerance levels $(\epsilon_{g})_{g\geq 0}$ that decrease gradually as $\infty=\epsilon_{0}>\epsilon_{1}>\epsilon_{2}>\ldots>0$. Each intermediate distribution (called a generation and denoted by $g$) is represented by a cloud of weighted particles$\left(\theta_{j},w_{j}^{g}\right)_{j=1,\ldots,J}$. We approximate each intermediate posterior distribution using a multivariate kernel density estimator (kde) denoted by $\pi_{\epsilon_{g}}(\theta|\mathcal{D})$. The parameters of the algorithm are the number of generations $G$, the population size $J$ (the number of particles in the cloud), and the number of Monte Carlo replications $R$ of $X$.
@@
where $e_{i}$ is an error term that captures the mismatch between the true value of the pure premium and its empirical counterpart estimated by the competitor insurance company using its historical data and the company-specific loading function. We further compare the observed commercial premiums to the model-generated ones via the distance defined in Section 3.2 with
@@
$w_{j}^{g}\propto\frac{\pi(\theta^{g}_{j})}{\pi_{\epsilon_{g-1}}(\theta)}\mathbb{I}_{d_{j}^{g}<\epsilon_{g-1}},\quad j=1,\ldots,J.$
The tolerance threshold is chosen so as to maintain a specified effective sample size (ess) of $J/2$ as in Del Moral et al. [8]. Following Kong et al. [14], the ess is estimated by $1/\sum_{j=1}^{J}(w_{j}^{g})^{2}$. This weighted sample then allows us to update the intermediate distribution as
$\pi_{\epsilon_{g}}(\theta|\mathcal{D})=\sum_{j=1}^{J}w_{j}^{g}K_{H}(\theta-\theta^{g}_{j}),$
where $K_{H}$ is a multivariate kde with smoothing matrix $H$. A common choice for the kde is the multivariate Gaussian kernel with a smoothing matrix set to twice the empirical covariance matrix of the cloud of particles $\{\theta_{j}^{g},w_{j}^{g}\}$ as in Beaumont et al. [4]. The procedure is summarized in Algorithm 1.
@@
12:    find $\epsilon_{g}\leq\epsilon_{g-1}$ so that $\widehat{\text{{ess}\@}}=\Bigl{[}\sum_{j=1}^{J}\left(w_{j}^{g}\right)^{2}\Bigr{]}^{-1}\approx J/2$, where
@@
13:    compute $\pi_{\epsilon_{g}}(\theta\mid\mathcal{D})=\sum_{j=1}^{J}w_{j}^{g}K_{\bm{H}}(\theta-\theta_{j}^{g})$
@@
The user must configure several aspects of the algorithm. The prior assumptions $\pi(\theta)$ determine the parameter space that will be searched. The loss ratio corridor $[\text{LR}_{\text{low}},\text{LR}_{\text{high}}]$ sets up the two regularization terms, ensuring that parameters associated with unreasonable pure premiums are excluded. The prior settings and loss ratio corridor can be guided by expert opinions. The population size $J$ drives the quality of the posterior distributions approximations through the cloud of particles. A large $J$ also enhances the chances of finding global optimums, as more particles improve the coverage of the parameter space. A greater number $R$ of Monte Carlo simulations ensures the accuracy of the pure premium evaluation. Both $R$ and $J$ contribute to the stability of the algorithm’s results over several runs. The number of generations $G$ relates to the tolerance level $\epsilon$, which in turn drives the narrowness of the posterior distribution output by the ABC algorithm. As one would expect, the computational time for the algorithm increases with higher values of $R,J$ and $G$. Therefore, the choice of suitable values for $G$, $J$, and $R$ can be made in consideration of a predetermined computational time budget. A practical solution to set $G$ is to stop the algorithm whenever the difference between two consecutive tolerance levels is lower than some threshold $\Delta_{\epsilon}$or if we reach a minimum tolerance level $\epsilon_{\min}$. Convergence results for ABC algorithms are readily available in the litterature as discussed in Remark 3.3.
@@
$\pi_{\epsilon_{G}}(\theta|\mathcal{D})=\sum_{j=1}^{J}w_{j}^{G}K_{H}(\theta-\theta^{G}_{j}),$ (9)
@@
Regarding convergence (2), as $\epsilon_{G}$ decreases toward 00, our ABC estimator converges to the distribution of $\theta$ conditional on $\mathcal{D}$. This is equivalent to the true posterior distribution only if $\mathcal{D}$ consists of sufficient statistics for the risk model $X$. Rubio and Johansen [19, Proposition 2] justify the use of non-sufficient statistics, under certain conditions, to make inferences about $\theta$. Their results remain valid when using a kernel density estimator such as (9) to represent the posterior distribution. The bias of the ABC posterior has been estimated to be $\mathcal{O}(\epsilon_{G}^{2})$ in Barber et al. [2].
@@
We illustrate the posterior distribution evolution along the algorithm iterations in 6.
@@
$J=1000\text{, }R=1000\text{, }\Delta_{\epsilon}=0.1\text{, and }LR\in[0.3,0.66].$
The algorithm halts at the $9^{\text{th}}$ generation reaching a tolerance level of $\epsilon=0.87$. Figure 6 shows the sequence intermediate posterior distributions for $\lambda$ and $\sigma$.
@@ 4 Methodology Assessment via Simulation
In this section, we embark on an empirical exploration, seeking to understand how the posterior distribution of the parameters behaves as the sample size $n$ increases. This experimentation has been designed to resemble as much as possible the real data situation considered in Section 5.3. We consider the risk, within a particular risk class, to be distributed as the random variable
$X=\sum_{k=1}^{N}U_{k},$
@@
The $U_{i}$’s are iid and independent from $N$. We suppose that we know the variance parameter $\sigma$ and we try to draw inference on $\lambda$ and $\mu$. The parameter values of the claim frequency and severity in (10) and (11) respectively are those infered in Section 5.3 for the $\mathsf{Poisson}-\mathsf{LogNorm}$ model using the mode estimator. The prior distributions are set to independent uniforms for $\lambda$ and $\mu$ as
@@
where the premium parameters $r$, $d$ and $l$ are sampled from that of the real data considered in Section 5, so that the simulated data is as close as possible to the real data. The $\eta_{i}$’s are iid from $\eta_{i}\sim\mathsf{Unif}([1.43,2.5])$, which corresponds to loss ratios between $40\%$ and $70\%$. We further set $\text{LR}_{\text{low}}=40\%$ and $\text{LR}_{\text{high}}=70\%$. We consider sample of sizes $25,50,100,$ and $200$. We configure the algorithm with a population size of $J=1,000$ and use $R=2,000$ Monte Carlo replications. To ensure the algorithm’s efficiency, we set a stopping threshold, requiring that the difference between two consecutive tolerance levels is smaller than $\Delta_{\epsilon}=1$ for the algorithm to halt. These settings are kept for the analysis of real-world data, as they strike a balanced compromise between accuracy and computing time. We generate $100$ samples of fake data and apply our procedure. Our goal is to compare the result obtained using our two pointwise estimators: the mean a posteriori map and the mode a posteriori mode. The estimators of the parameters $\lambda$ and $\mu$ are given on Figure 7.
@@
$\overline{\text{LR}}=\frac{1}{n}\sum_{i=1}^{n}\frac{p_{i}}{\tilde{p}_{i}}.$
@@ 5.1 Evolution and growth of the pet insurance market
To date, the adoption and acceptance of pet insurance still varies significantly across regions of the world. Nordic countries, such as Sweden, have historically had a very high penetration rate with around $70\%$ of pets insured. Some Anglo-Saxon countries (UK and Germany mostly) have seen significant growth in the pet insurance market during the last decades, leading to $30\%$ of penetration rate. Other developed countries, like France, have significantly lower market sizes, with less than $10\%$ of pets that are insured, which suggests high growth potential. The market place for pet insurance in the USA is currently also experiencing sustained growth. According to MarketWatch nationwide survey [^6] https://www.marketwatch.com/guides/pet-insurance/pet-insurance-facts-and-statistics/, about $44.6\%$, of pet owners stated they currently have pet insurance. The North American Pet Health Insurance Association (NAPHIA) conducted a survey for its 2022 State of the Industry Report and found that over $4.41$ million pets were insured in North America in 2021, up from $3.45$ million in 2020. The report also revealed that pet insurance premiums totaled \$$\$2.84$ billion in 2021, marking a $30.5\%$ increase from the previous year.
@@ 5.2 Data description
[…]
@@
The first five rows are given in Table 2.
@@
We have collected rates associated to $12$ risk classes given in Table 3.
Risk class ##\## specie breed gender age
@@
Remark 5.1 briefly discusses data quality.
@@
We conduct two separate studies. In Section 5.3, we focus on a specific risk class associated to a female, 4 years old, australian sheperd. Several claim models are compared. One is selected to look into the pricing strategies of the actors. In Section 5.4, we look into the quotes of various risk classes that we investigate using a single model.
@@ 5.3 Analysis of one risk class using several models
Combining the distributions for the claim frequency and severities results in a total of $3$ loss models. The population size in the abc algorithm is set to $J=1{,}000$. The pure premiums are computed using $R=2{,}000$ Monte Carlo replications. The algorithm stops whenever the difference between two consecutive tolerance levels is lower than $\Delta_{\epsilon}=1$. The bounds for the loss ratio corridor are set to $\text{LR}_{\text{low}}=40\%$ and $\text{LR}_{\text{high}}=70\%$. The posterior distributions of the parameters for each model are provided in Figure 10.
@@
For all the models, the algorithm updates the prior distribution in an informative way. Table 4 provides the tolerance levels (ranked in increasing order) during the last iteration of the abc algorithm for the loss models.
@@
The final tolerance levels lies between $93.72$ and $113.15$ which is higher than the tolerance obtained in the simulation study which was around $33$ for $50$ data points and $50$ for $200$ data points. This discrepancy indicates misspecifications which stem from our assumptions about insurance companies adhering to the expectation principle for premium calculation and the models employed for claim frequency and claim amounts. Table 5 reports the estimations of the parameters of all the model using the map and the mode.
@@
Table 6 reports the estimations of the average total claim amounts and the average loss ratio for all the models for all models when fitted using the map and the mode.
@@
We note that the risk level, characterized here by the expected total claim amount, is similar for all the models, maybe a bit higher for the model having the $\mathsf{LogNorm}(\mu=0,\sigma)$ as claim sizes distribution. We further look into the loading function approximated via the isotonic regression. We estimate the pure premium for each model using the map as an estimator of the model parameters and we plot the isotonic regression function to explain the commercial premium on Figure 11.
@@
The isotonic fits of the loading function accross all the models are similar which means that the models all agree on a common ordering of the pure premiums of the various insurance coverages. To highlight the explanatory power of our methodology, let’s focus on the $\mathsf{Poisson}(\lambda)-\mathsf{LogNorm}(\mu,\sigma=1)$ loss model. Note that the choice of the loss model is somewhat arbitrary because the information extracted from the data in Figure 11 is relatively consistent across the considered models. In Figure 12, we present a plot that illustrates the relationship between the commercial premium and the pure premium for the $\mathsf{Poisson}(\lambda)-\mathsf{LogNorm}(\mu,\sigma=1)$ model. Different insurance companies are indicated by distinct colors, providing a visual representation of each company’s respective rates.
@@ 5.4 Analysis of several risk classes with one model
The $\mathsf{Poisson}(\lambda)-\mathsf{LogNorm}(\mu,\sigma=1)$ model is fitted to the data within each risk classes (90 quotes) of Table 3. The prior settings are given by
@@
$J=1000,\text{ }R=2000,\text{ and }\Delta_{\epsilon}=1.$
Figure 13 shows the posterior predictive distribution of the expected total claim amounts and the averaged loss ratio within each risk class.
@@
Figure 13a allows us to compare the different risk classes. We note that an older dog is more expensive on average and that the breeds may be ordered as Australian Sheperd, Golden-Retriever, German sheperd and french bulldog in terms of riskiness. Figure 13b indicates that the loss ratios are arround $62-65\%$ for all the risk classes.

### v2

@@ 3.3 Population Monte Carlo Approximate Bayesian Computation algorithm
[…]
@@
where $e_{i}$ is an error term that captures the mismatch between the true value of the pure premium and its empirical counterpart estimated by the competitor insurance company using its historical data and the company-specific loading function. We further compare the observed commercial premiums to the model-generated ones via the distance defined in Section˜3.2 with
@@
$w_{j}^{g}\propto\frac{\pi(\theta^{g}_{j})}{\pi_{\epsilon_{g-1}}(\theta)}\mathbb{I}_{d_{j}^{g}<\epsilon_{g}},\quad j=1,\ldots,J.$
The tolerance threshold is chosen so as to maintain a specified effective sample size (ess) of $J/2$ as in Del Moral et al. [8]. Following Kong et al. [14], the ess is estimated by $1/\sumop\displaylimits_{j=1}^{J}(w_{j}^{g})^{2}$. This weighted sample then allows us to update the intermediate distribution as
$\pi_{\epsilon_{g}}(\theta|\mathcal{D})=\sumop\displaylimits_{j=1}^{J}w_{j}^{g}K_{H}(\theta-\theta^{g}_{j}),$
where $K_{H}$ is a multivariate kde with smoothing matrix $H$. A common choice for the kde is the multivariate Gaussian kernel with a smoothing matrix set to twice the empirical covariance matrix of the cloud of particles $\{\theta_{j}^{g},w_{j}^{g}\}$ as in Beaumont et al. [4]. The procedure is summarized in Algorithm˜1.
@@
12:  find $\epsilon_{g}\leq\epsilon_{g-1}$ so that $\widehat{\text{{ess}{}}}=\Bigl[\sumop\displaylimits_{j=1}^{J}\left(w_{j}^{g}\right)^{2}\Bigr]^{-1}\approx J/2$, where
@@
13:  compute $\pi_{\epsilon_{g}}(\theta\mid\mathcal{D})=\sumop\displaylimits_{j=1}^{J}w_{j}^{g}K_{\bm{H}}(\theta-\theta_{j}^{g})$
@@
The user must configure several aspects of the algorithm. The prior assumptions $\pi(\theta)$ determine the parameter space that will be searched. The loss ratio corridor $\text{LR}_{\text{low}},\text{LR}_{\text{high}}]$ sets up the two regularization terms, ensuring that parameters associated with unreasonable pure premiums are excluded. The prior settings and loss ratio corridor can be guided by expert opinions. The population size $J$ drives the quality of the posterior distributions approximations through the cloud of particles. A large $J$ also enhances the chances of finding global optimums, as more particles improve the coverage of the parameter space. A greater number $R$ of Monte Carlo simulations ensures the accuracy of the pure premium evaluation. Both $R$ and $J$ contribute to the stability of the algorithm’s results over several runs. The number of generations $G$ relates to the tolerance level $\epsilon$, which in turn drives the narrowness of the posterior distribution output by the ABC algorithm. As one would expect, the computational time for the algorithm increases with higher values of $R,J$ and $G$. Therefore, the choice of suitable values for $G$, $J$, and $R$ can be made in consideration of a predetermined computational time budget. A practical solution to set $G$ is to stop the algorithm whenever the difference between two consecutive tolerance levels is lower than some threshold ϵor if we reach a minimum tolerance level $\epsilon_{\min}$. Convergence results for ABC algorithms are readily available in the litterature as discussed in Remark˜3.3.
@@
$\pi_{\epsilon_{G}}(\theta|\mathcal{D})=\sumop\displaylimits_{j=1}^{J}w_{j}^{G}K_{H}(\theta-\theta^{G}_{j}),$ (9)
@@
Regarding convergence (2), as $\epsilon_{G}$ decreases toward 0, our ABC estimator converges to the distribution of $\theta$ conditional on $\mathcal{D}$. This is equivalent to the true posterior distribution only if $\mathcal{D}$ consists of sufficient statistics for the risk model $X$. Rubio and Johansen [19, Proposition 2] justify the use of non-sufficient statistics, under certain conditions, to make inferences about $\theta$. Their results remain valid when using a kernel density estimator such as (9) to represent the posterior distribution. The bias of the ABC posterior has been estimated to be $\mathcal{O}(\epsilon_{G}^{2})$ in Barber et al. [2].
@@
We illustrate the posterior distribution evolution along the algorithm iterations in ˜6.
@@
$J=1000\text{, }R=1000\text{, }{}_{\epsilon}=0.1\text{, and }LR\in[0.3,0.66].$
The algorithm halts at the $9^{\text{th}}$ generation reaching a tolerance level of $\epsilon=0.87$. Figure˜6 shows the sequence intermediate posterior distributions for $\lambda$ and $\sigma$.
@@ 4 Methodology Assessment via Simulation
In this section, we embark on an empirical exploration, seeking to understand how the posterior distribution of the parameters behaves as the sample size $n$ increases. This experimentation has been designed to resemble as much as possible the real data situation considered in Section˜5.3. We consider the risk, within a particular risk class, to be distributed as the random variable
$X=\sumop\displaylimits_{k=1}^{N}U_{k},$
@@
The $U_{i}$’s are iid and independent from $N$. We suppose that we know the variance parameter $\sigma$ and we try to draw inference on $\lambda$ and $\mu$. The parameter values of the claim frequency and severity in (10) and (11) respectively are those infered in Section˜5.3 for the $\mathsf{Poisson}-\mathsf{LogNorm}$ model using the mode estimator. The prior distributions are set to independent uniforms for $\lambda$ and $\mu$ as
@@
where the premium parameters $r$, $d$ and $l$ are sampled from that of the real data considered in Section˜5, so that the simulated data is as close as possible to the real data. The $\eta_{i}$’s are iid from $\eta_{i}\sim\mathsf{Unif}([1.43,2.5])$, which corresponds to loss ratios between $40\%$ and $70\%$. We further set $\text{LR}_{\text{low}}=40\%$ and $\text{LR}_{\text{high}}=70\%$. We consider sample of sizes $25,50,100,$ and $200$. We configure the algorithm with a population size of $J=1,000$ and use $R=2,000$ Monte Carlo replications. To ensure the algorithm’s efficiency, we set a stopping threshold, requiring that the difference between two consecutive tolerance levels is smaller than $_{\epsilon}=1$ for the algorithm to halt. These settings are kept for the analysis of real-world data, as they strike a balanced compromise between accuracy and computing time. We generate $100$ samples of fake data and apply our procedure. Our goal is to compare the result obtained using our two pointwise estimators: the mean a posteriori map and the mode a posteriori mode. The estimators of the parameters $\lambda$ and $\mu$ are given on Figure˜7.
@@
$\overline{\text{LR}}=\frac{1}{n}\sumop\displaylimits_{i=1}^{n}\frac{p_{i}}{\tilde{p}_{i}}.$
@@ 5.1 Evolution and growth of the pet insurance market
To date, the adoption and acceptance of pet insurance still varies significantly across regions of the world. Nordic countries, such as Sweden, have historically had a very high penetration rate with around $70\%$ of pets insured. Some Anglo-Saxon countries (UK and Germany mostly) have seen significant growth in the pet insurance market during the last decades, leading to $30\%$ of penetration rate. Other developed countries, like France, have significantly lower market sizes, with less than $10\%$ of pets that are insured, which suggests high growth potential. The market place for pet insurance in the USA is currently also experiencing sustained growth. According to MarketWatch nationwide survey [^6] https://www.marketwatch.com/guides/pet-insurance/pet-insurance-facts-and-statistics/, about $44.6\%$, of pet owners stated they currently have pet insurance. The North American Pet Health Insurance Association (NAPHIA) conducted a survey for its 2022 State of the Industry Report and found that over $4.41$ million pets were insured in North America in 2021, up from $3.45$ million in 2020. The report also revealed that pet insurance premiums totaled $\mathdollar 2.84$ billion in 2021, marking a $30.5\%$ increase from the previous year.
@@ 5.2 Data description
During the week of the $18\textsuperscript{th}$ of May 2024, we have collected $1,080$ quotes from $5$ insurance companies. Each row of our datasets corresponds to a yearly premium collected from some insurance company website associated to a specific insurance coverage and a specific dog. We therefore find the coverage parameters which are the coverage rate $r$, the deductible $d$ and the limit $l$. Recall that the compensation for an annual expense of amount $X$ is calculated as $\min\left[\max(r\cdot X-d,0),l\right]$. We also have the rating factors which reduces for pet insurance in France to specie, breed, age and gender. Table˜1 provides a list of the variables in the datasets.
@@
The first five rows are given in Table˜2.
@@
We have collected rates associated to $12$ risk classes given in Table˜3.
Risk class #\# specie breed gender age
@@
Remark˜5.1 briefly discusses data quality.
@@
We conduct two separate studies. In Section˜5.3, we focus on a specific risk class associated to a female, 4 years old, australian sheperd. Several claim models are compared. One is selected to look into the pricing strategies of the actors. In Section˜5.4, we look into the quotes of various risk classes that we investigate using a single model.
@@ 5.3 Analysis of one risk class using several models
Combining the distributions for the claim frequency and severities results in a total of $3$ loss models. The population size in the abc algorithm is set to J=1, [^0] J=1{,}000. The pure premiums are computed using R=2, [^0] R=2{,}000 Monte Carlo replications. The algorithm stops whenever the difference between two consecutive tolerance levels is lower than $_{\epsilon}=1$. The bounds for the loss ratio corridor are set to $\text{LR}_{\text{low}}=40\%$ and $\text{LR}_{\text{high}}=70\%$. The posterior distributions of the parameters for each model are provided in Figure˜10.
@@
For all the models, the algorithm updates the prior distribution in an informative way. Table˜4 provides the tolerance levels (ranked in increasing order) during the last iteration of the abc algorithm for the loss models.
@@
The final tolerance levels lies between $93.72$ and $113.15$ which is higher than the tolerance obtained in the simulation study which was around $33$ for $50$ data points and $50$ for $200$ data points. This discrepancy indicates misspecifications which stem from our assumptions about insurance companies adhering to the expectation principle for premium calculation and the models employed for claim frequency and claim amounts. Table˜5 reports the estimations of the parameters of all the model using the map and the mode.
@@
Table˜6 reports the estimations of the average total claim amounts and the average loss ratio for all the models for all models when fitted using the map and the mode.
@@
We note that the risk level, characterized here by the expected total claim amount, is similar for all the models, maybe a bit higher for the model having the $\mathsf{LogNorm}(\mu=0,\sigma)$ as claim sizes distribution. We further look into the loading function approximated via the isotonic regression. We estimate the pure premium for each model using the map as an estimator of the model parameters and we plot the isotonic regression function to explain the commercial premium on Figure˜11.
@@
The isotonic fits of the loading function accross all the models are similar which means that the models all agree on a common ordering of the pure premiums of the various insurance coverages. To highlight the explanatory power of our methodology, let’s focus on the $\mathsf{Poisson}(\lambda)-\mathsf{LogNorm}(\mu,\sigma=1)$ loss model. Note that the choice of the loss model is somewhat arbitrary because the information extracted from the data in Figure˜11 is relatively consistent across the considered models. In Figure˜12, we present a plot that illustrates the relationship between the commercial premium and the pure premium for the $\mathsf{Poisson}(\lambda)-\mathsf{LogNorm}(\mu,\sigma=1)$ model. Different insurance companies are indicated by distinct colors, providing a visual representation of each company’s respective rates.
@@ 5.4 Analysis of several risk classes with one model
The $\mathsf{Poisson}(\lambda)-\mathsf{LogNorm}(\mu,\sigma=1)$ model is fitted to the data within each risk classes (90 quotes) of Table˜3. The prior settings are given by
@@
$J=1000,\text{ }R=2000,\text{ and }{}_{\epsilon}=1.$
Figure˜13 shows the posterior predictive distribution of the expected total claim amounts and the averaged loss ratio within each risk class.
@@
Figure˜13a allows us to compare the different risk classes. We note that an older dog is more expensive on average and that the breeds may be ordered as Australian Sheperd, Golden-Retriever, German sheperd and french bulldog in terms of riskiness. Figure˜13b indicates that the loss ratios are arround $62-65\%$ for all the risk classes.